What is the angular velocity of a hour and minute hand of a clock?With Explanations plz best explanation.

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

The angular displacement of each of the hour and minute hand is2\pi radiansfor one complete rotation.

The time taken for this one complete rotation by the hour hand is 1 hour=3600s

Hence the angular velocity of hour hand is\dfrac{2\pi}{3600}rad/s=\dfrac{\pi}{1800}rad/s

Similary angular velocity of the minute hand is\dfrac{2\pi}{60}rad/s=\dfrac{\pi}{30}rad/s